How do I solve this?
if a + b = 25 what is 2a + 2b + 4

Grade:8

1 Answers

Abhinav Gupta
6 years ago
a+b =25. Multiply the equation by 2.. 2a+2b= 50. Therefore,. 2a+2b+4=50+4=54. So the answer is 54....